Project

General

Profile

Actions

Additional issues arising from Cycle 2017/01

Plucked from the eLog.
Issues already fixed are in italics.
Actions/assignments are in bold.

Post Cycle 2017/02 Startup discussion


Other

  • A/C Unit 3 alarm
    • What do we do when John's away - P Masterton, ISIS (need to chase units 3 and 4 - Colin?)
  • "On Run Status, the Integrated TOF1 (requested) Triggers and Requested Triggers numbers differ by 2: 76810 vs 76812."
    • On various runs we've seen differences of up to 10 triggers. Why? Need to check number of reported integrated trips matches number of triggers in file: Action DR
  • "U/S Tracker - LED system isn't working; possibly a switch got flipped." Check during run-up: MUchida and PKyberd
  • Tracker volume flush read-back broken - add to checklist for run-up : Action for SB Start Helium flush next week; check read-back in test runs: MUchida
  • MOM checklist should include checking that the State Machines are in the appropriate states. (which states are those?) AKurup will write a procedure on how to restart SM's, ALH and then mask out known false alarms Action AKurup to produce a list of the states appropriate for running for MOM to check ; MOM to work through it at start of running ; Action SBoyd to modify MOM checklist
  • "EPICS reports that A/C unit 2 - which I believe is actually the reading from unit 3 - is reporting 26.3°C suggesting that the East end of the Hall is getting warm"
    • We need to understand the Hall temperatures: numbers presently range from 22 to 27 °C: Range of 5°C is not huge. Check with D.C. which are the sensitive items. Sensitive equipment should have local monitoring anyway.

"We have selected "LH2 Empty" in Run Control: it STILL wrote "LH2" into the CDB! CDB viewer's reporting "LH2" as the material and "Empty" as the shape which I think is the wrong way round - not sure if this is the viewer or Run Control." Need to check CDB viewer displays correctly (DR/JM) - it's correct ; Need to update for Neon (which is what we actually had) - stands DR

  • "discovered during this that the overtemperature alarms had been LEFT SILENCED at some point, without this being passed on to the next shift."
    Shifters silencing alarms rather than acknowledging, and/or not expanding the tree to identify the individual alarm MOM/SteveBoyd
    • e.g. If you silence the Proton Absorber Status because it alarms when changing Beamline settings, how will you know if it droops open during the run (c.f. December)?
    • Status of Shift Operation Checklist? (SBoyd) Improve via the checklist: Action on SB to add alarm check to shift changeover checklist
  • "MOM/DC Configuration Checklist" omitted add to MOM/DC handover Action SB

Controls and Monitoring

  • Calibration of compressed air pressure sensors - done
  • Readouts from A/C units 2 and 3 swapped
  • Massive number of bogus alarms (e.g. on a PV that is inValid):
    • "Tracker 1 Cryostat - vacuum gauge - alarm as reading is inValid" - Tracker alarms need to be re-enabled; MUchida to notify DC when ready, probably next week.
    • "I have disabled the "QPS Signals", "QPS Interlocks" and "Interlocks" trees on the SSU Alarm Handler, to match the SSD Alarm Handler State." - should be covered in the SM fix. AK needs to look at removing the alarms on the QPS; SSU should generally match SSD; fix via spreadsheets.
    • "I've also disabled the alarm on the Neutron Monitor - which is no longer connected so gives a permanent error." - Taken out completely
    • "AC5 status disabled in ENV alarm handler" - Unit will not be networked, PVs to be removed from spreadsheet
    • "Henry has disabled the following in the alarm handler in consultation with MOM-Ed:
      • TKU: GHe, Cryo1 vacuum
      • TKD: GHe, Cryo1/2 (really 3/4) vacuum, Cryo2-LCRB-temp2 (actually Cryo4LC-temp2), Cryo2-LCRB-AFEheater2 (also mislabelled)"
        GHe monitoring boxes restarted; vacuum gauges replaced and the readout isn't working (AK/CMw); temp and AFE: need to check again; mislabelling fixed
    • "I've disabled the Q4 and Q7 MPS water flows in the Alarm Handler because of the false alarms ... I am suspicious of the fact that we only get these false alarms on those two supplies" - Comes from the readback glitches - Time filter settings have been re-applied and should prevent the alarms HN to check present alarm status
    • FCD Load Cells - see below - HW disconnect - not a C&M thing
    • What is the plan for SSU CC1 and 2? see below - Need to check cable - comms issue between compressor and IOC
  • "the limits on SSU CC3 High-Side Pressure have been adjusted to get rid of the frequent alarms we were getting a couple of days ago. (now just to get CC1 and 2 sorted out)."
    • What is the plan for SSU CC1 and 2? - Need to check cable - comms issue between compressor and IOC
  • "SSU Alarm Handler DISABLED in all channels" - should be covered in the SM fix
  • "Run Control is hung up; restarted in separate terminal" - Run Control has long-standing deep-seated issues: Unfixable. Use it slowly.
  • "Will also check ... reported water leak in trench. "
    "About 13:00 Ajit restarted the Environment IOC - this should have got the air temperature and other sensors going."
    • SecurityProbe system not being read out by the IOC - needs some sort of heartbeat alarm: check the Read-back Status of the PV via Nagios: action PF/AK ; Stale info from SecurityProbe system has safety implications
  • "Noticed power supply for EMR Single Anode PMT # 26 was tripped off and probably had been for a while."
    "Noticed also that in the Detectors ALH the entire EMR tree is Disabled, although Ed did put the Detectors into Running yesterday." - should be covered in the SM fix
    • Not clear when the ALH EMR tree was turned off. Who is responsible for checking the Alarm Handlers over before running: experts or C&M? - see below
    • MOM checklist should include checking that the State Machines are in the appropriate states. (which states are those?) AKurup will write a procedure on how to restart SM's, ALH and then mask out known false alarms ; MOM to work through it at start of running ; MOM to hold the list of known issues. ; SBoyd to oversee MOMs
  • "Roof water temperature caused a red alarm, upper limit. - Returned to normal before a cause to be found. Assumed to be a measurement fluctuation."
    • We see glitches in the roof water PV MICE-SER-H2O-01:TPrimary: 6:06 31 May 2017 Checked Archiver - glitch in output variable isn't accompanied by any glitch in its inputs. SBoyd: ask shifters to keep an eye on water (stripcharts?).
  • "We have selected "LH2 Empty" in Run Control: it STILL wrote "LH2" into the CDB! CDB viewer's reporting "LH2" as the material and "Empty" as the shape which I think is the wrong way round - not sure if this is the viewer or Run Control." Need to check what is being written out (and displayed) by Run Control (AK) (done)
  • "I just realised that one reason the Alarm Handler is such a pain to use is that the software is incapable of scrolling and making a beep noise at the same time!" Test if scrolling and beeping works on miceiocpctest - VNC? _On miceiocpctest you cannot scroll while is beeping (PF). Can't fix.
  • "There was a major alarm from DATE at about 1360 spills/ 83k triggers, but by the time we found it we couldn't see any cause or error message. No sign of a trip in the beamloss strip chart. The transfer of the ISIS beamloss numbers from ISIS into EPICS is still very unpredictable - sometimes we get the values updated on 2 successive spills, sometimes the values are stay frozen for 9 or 10 spills." - Can't see a 10-spill plateau in the Archiver, so we could lose an ISIS trip from our own monitoring. needs beam
  • Repeating alarms need thresholds adjusting putting them on a strip-chart makes it easier to see if there's a real problem AKurup will update via spreadsheets. Then look out for known alarms during start-up.
    • "SSU CC3 High-Side Pressure"
    • DS PSU voltage Check with Mike
    • "Q4 and Q7 are showing a consistent warning on the MPS water flow" #1880
    • "alarm from Q2 deltaT" #1880
    • "Q8 reports temperature change warning alarms." #1880
    • "Q8 has a warning on the voltage" #1880
    • "Supply air pressure reach 6.8 Bar before returning to normal." - increase Limit Hi 7.0, HiHi 7.2
    • "Alarms in beam line: quad delta T touches limit" : #1880
  • "Heartbeat alarms for miceecserv1 and miceraid 5. The help pop-ups flash up an illegible error message." - Now Fixed
    • In Alarm Handlers, we need working pop-ups ...
    • Changed during run?
  • "After acknowledging the miceraid5 alarm, the PV remains shown as red, but the alarm keeps returning. This seems to be because the Heartbeat re-sets the status to zero before making the next test, which fails again. Probably, this is Wrong as it operates differently to the majority of other alarms."
    • Is the Heartbeat test operation correct? AK/Paolo to see if it can be made consistent Needs also to then respond to later Nagios alerts on the same server. Maybe look at time stamp.
  • "temperature warnings from the CS1 VME crate."
    • Need to clarify temperatures of crates - are they really warmer than expected, else tweak limits AK to track down appropriate expert. ; There is an action on getting the crates cleaned out. Refer action to D.C. or Craig, recommend clean out crates before run.

Beamline

  • Beamstop circuit 1 leak - tested, not failed this time
  • Beamstop Restraint possibly has stripped bolt - bolts replaced, updated documentation passed to Steve Boyd to put on wiki , Need to arrange BLOC refreshers, SB to check documentation is posted
  • "We have pulled out the target one notch to 35.75 mm BCD, as the losses are now much spikier since the work ISIS did this morning"
    • Beamline Co-ordinator has raised with ISIS... HN to remind Bryan
  • DS Shield Pressures DS team cleaning gas over summer
  • "Decay solenoid supply pressure PI21 - fluctuated back to normal"
    • DS needs gas cleaning and TLC DS team cleaning gas over summer
  • DS PSU voltage alarm : Check with Mike widen limits
  • Tweak MPS temperatures difference limits: #1880
  • Target laser saga, and control PC / R78 - saga continues. Don't expect laser readout into EPICS this cycle.
  • Question from ISIS about the MICE 3ms spill gate - why not 1 ms?
    • Possible balance between inefficiency from dead-time vs. activation of ring for no data
    • How are the good muons distributed within the MICE spill? Won't be answered before cycle.

Updated by Nebrensky, Henry over 6 years ago · 43 revisions